OpenStudy (anonymous):

Factor 80y^2-120y+45

OpenStudy (anonymous):

u can take common first and then use quadratic formula..

OpenStudy (anonymous):

5(80y^2-120y+45) 16y^2-24y+9 \[24\pm \sqrt{576-576}/32\]

OpenStudy (anonymous):

=24/32 =3/4